A real refrigerator works between two temperature reservoirs, one at 277 Kelvin, and one at 345 Kelvin. In order to maintain a constant heat flow, during one cycle 350 Joules of heat is transferred from the cold reservoir, while the net work in one cycle is 500 Joules.
(a) Obtain the coefficient of performance for this refrigerator.
(b) Compare the result in (a) with the ideal case. That is, what is the maximum coefficient of performance if this was an ideal refrigerator.

please help asap thanks!
Answer #1

Tc = 277 K , Th = 345 K, Qc = 350 J . W = 500 J

a) coefficient of performance for refrigerator is

COP = Qc/W = 350 /500

COP = 0.7

b) In ideal case

COP = Tc/(Th -Tc )

= 277/(345-277)

Maximum COP =4.07

COP of refrigerator is less than COP in ideal case.